A car drives 500 ft on a road that is inclined 12° to the horizontal. The car weighs 2500 lb. Thus gravity acts straight down on the car with a constant force F = -2500j. Find the work done by
the car in overcoming gravity.

Answer by ikleyn(52786)

.

Due to the  Conservation  Law of mechanical energy,  the work done by the car engine
is equal to the change of the potential energy of the car.

The change of the potential energy of the car is

        2500*500*sin(12°) = 2500*500*0.2079 = 259875 lb*ft, approximately.         ANSWER

Solved, answered and explained.